2-49) Randy Duckout has been asked to develop an esti- mate of the per-unit selling price (the price that each unit will be sold for) of a new line of handcrafted booklets that offer excuses for missed appoint- ments. His assistant Doc Duckout has collected information that Randy will need in developing his estimate: Cost of direct labor Cost of materials $20 per hour $43.75 per batch of 25 booklets Cost of overhead items 50% of direct labor cost Desired profit 20% of total manufacturing cost Doc also finds out that (1) they should use a 75% learning curve for estimating the cost of direct labor, (2) the time to complete the 1st booklet is estimated at 0.60 hour, and (3) the estimated time to complete the 25th booklet should be used as their standard time for the purpose of determining the unit selling price. What would Randy and Docs estimate be for the unit selling price?
Answer #1

from 75% learning total time taken=0.09*25=2.25
total manufacturing cost=20*2.25+43.75+0.5*20*2.25=$74.125
total cost=$74.125*1.2=$88.95
per unit cost=$3.558